# TRIM: Achieving Extreme Sparsity with Targeted Row-wise Iterative Metric-driven Pruning

### v1

@@ 4.1 Main results
Although we primarily evaluate TRIM with the Wanda (Sun et al., 2024) metric, the dimension-wise sparsity and per-output optimization of TRIM can be applied to other pruning metrics. To test the generalization capabilities of TRIM, we evaluate using the SparseGPT pruning metric (Frantar and Alistarh, 2023), which assigns a score to each weight based on the squared value of the weight divided by the corresponding diagonal element of the inverse Hessian approximation $((\mathbf{X}^{T}\mathbf{X})^{-1})_{ii}$, where $\mathbf{X}$ are the input activations to the layer. SparseGPT groups weights in blocks of 128 inputs (columns of the weight matrix), as opposed to the per-output (a single row) comparison group of Wanda. We omit the weight updates in SparseGPT to focus on the pruning configuration and the potential improvement capabilities of TRIM. We also evaluate the magnitude pruning metric (Han et al., 2015), which uses the absolute value of the weights as a score, $|\mathbf{W_{ij}}|$, and selects pruning candidates by comparing weights within the entire layer. Table 3 shows that the per-output pruning configuration with TRIM performs best, confirming that TRIM generalizes across different metrics.
Table 3: Wikitext validation perplexity of LLaMA-2-13B and Qwen2.5-14B pruned at 70% sparsity with OWL layer ratios. TRIM improves upon the per-output configuration and generalizes to other pruning metrics besides Wanda. Underscored results indicate the default pruning configuration.
Model Method Pruning Metric (layer) (input, 128) (1, output) (1, output) +TRIM LLaMA-2-13B Magnitude $|\mathbf{W_{ij}}|$ 59.86 49.25 23.48 19.94 Wanda $|\mathbf{W_{ij}}|\cdot||\mathbf{X_{j}}||$ 20.36 24.64 16.02 14.39 SparseGPT $[|\mathbf{W}|^{2}/\text{diag}(\mathbf{H^{-1}})]_{ij}$ 19.75 19.23 15.19 14.11 Qwen2.5-14B Magnitude $|\mathbf{W_{ij}}|$ 2e3 2e3 3e4 2e3 Wanda $|\mathbf{W_{ij}}|\cdot||\mathbf{X_{j}}||$ 40.92 43.54 36.46 33.21 SparseGPT $[|\mathbf{W}|^{2}/\text{diag}(\mathbf{H^{-1}})]_{ij}$ 24.71 27.93 24.18 23.50
@@ 5 Analysis
An output dimension’s pruning sensitivity is closely tied to its importance score concentration. We examine the concentration of Wanda scores ($|W_{i,:}|\cdot||X||_{2}$) for each output dimension using Gini coefficients. Figure 3 displays histograms of these per-dimension Gini coefficients for the first, middle and last layer of Qwen2.5-14B. Figure 3 demonstrates that some dimensions rely on a few key weights while others distribute importance more broadly. These trends are consistent across all models and model layers.
(a) Layer 1
(b) Layer 24
(c) Layer 48
Figure 3: Histograms of Gini coefficients for the Wanda pruning metric across output dimensions, in different layers of Qwen2.5-14B (K-proj). A high Gini coefficient indicates that the signal for that dimension is concentrated in a few weights.
The heterogeneity of importance concentration translates directly into differing pruning sensitivities across output dimensions. To demonstrate this, we increase the uniform sparsity ratio applied to a layer and measure the resulting quality degradation for each individual output dimension. We quantify the quality using the cosine similarity $c_{i}=\text{CosineSimilarity}(Y_{i,:},\hat{Y}_{i,:})$ between the original output $Y_{i,:}$ and the pruned output $\hat{Y}_{i,:}$ for that dimension, given a set of sample inputs $X$. Figure 4 plots these cosine similarities $c_{i}$ against increasing levels of sparsity $T$. The rate at which quality degrades differs significantly across dimensions. Some dimensions maintain high similarity even at high sparsity levels (indicating resilience to pruning), while others experience a rapid drop even at lower sparsity (indicating sensitivity to pruning). This result confirms that output dimensions are indeed differently “difficult” to prune and that this difficulty is linked to their importance concentration.
Figure 4: Cosine similarity for all output dimensions (Qwen2.5-14B, layer-24, k-proj). Increasing sparsity shows varying pruning sensitivity.
@@
Pruning can remove redundant weights with minimal impact at low sparsity levels due to the overparameterization of LLMs. However, as sparsity rises, pruning inevitably affects weights carrying more essential activation signals, leading to a faster decay in performance. This non-linear decline is evident in the average trend of quality degradation on the dimension-level (Figure 4) or in downstream metrics like the exponential increase in perplexity at high sparsity rates (Figure 2).
@@
Beyond the differences in their sensitivity to pruning, output dimensions vary significantly in their importance for the model’s performance. To illustrate this, we conduct an experiment with Qwen2.5-3B, where we fully eliminate exactly one output dimension per layer, resulting in 0.03% total sparsity (Table 5). Removing the dimension with the smallest L2-norm only increases perplexity by 0.16 from the unpruned baseline, similar to choosing a random dimension (+0.15). However, pruning the dimension with the highest norm increases perplexity to 273.10. The importance of a dimension can also be linked to the concentration of outlier positions present in it. We define an outlier position using the OWL criterion, where the Wanda metric score $A_{ij}$ is at least a multiple of the layer mean $\bar{A}$. We take the top 10% of dimensions with the largest number of outliers (i.e., “outlier-dense” dimensions) and prune them with 90% sparsity, resulting in 9% total model sparsity. Note that pruning this way retains all of the outlier positions since they are always the last weights to get pruned. Table 5 shows that pruning outlier-dense dimensions significantly impacts performance compared to randomly pruning an equivalent number of dimensions with the same sparsity. This result highlights two key points.
Method Perplexity
Baseline 8.03
$D$ w. min norm 8.19
$D$ w. max norm 273.10
Random $D$ 8.18
Table 5: Wikitext validation perplexity. 10% of dimensions $D$ are heavily pruned (90%). Outlier-dense dimensions are important, however differently so across model families.
Model Baseline Random Outlier-dense $D$
Qwen2.5-7B 6.85 9.17 26.39
Qwen2.5-14B 5.29 7.70 26.16
LLaMA2-7B 5.47 7.27 255.06
LLaMA2-13B 4.88 6.23 74.27
First, outlier-dense dimensions are important for model performance. Second, although outlier positions in an outlier-dense dimension may appear to capture most of the signal, our findings show that non-outlier weights also carry significant informational value, limiting the effectiveness of aggressive pruning. Further, the extent of this impact differs between model families. For example, LLaMA models exhibit a more pronounced degradation than Qwen, indicating that outlier-dense dimensions are more important for some models than for others. Together, these findings highlight that output dimensions are not uniformly important and that simple heuristics, such as L2-Norm or outlier characteristics, can identify dimensions critical for preserving model capabilities.
@@ 6 Discussion
Targeted sparsity allocation has proven successful for enhancing the performance of pruned LLMs, but has so far only been attempted at the layer level (Yin et al., 2024; Lu et al., 2024; Sun et al., 2025; Xu et al., 2025). This work extends this paradigm to individual output dimensions within each layer. This is particularly promising, and indeed necessary, when aiming for high sparsity levels. As our findings indicate (Observation 5), post-pruning quality tends to degrade at an accelerated rate with increasing sparsity, making indiscriminate or overly coarse pruning strategies increasingly detrimental. The core premise of TRIM and the concept of dimension-wise sparsity allocation is that there exists significant heterogeneity to exploit at the dimension level (Observation 5 and 5). However, moving to such a fine-grained level of control introduces challenges: the search space for optimal sparsity configurations becomes substantially larger, and the feedback signals (e.g., quality metrics per dimension) might become less reliable or more noisy compared to layer-level aggregates.
Figure 5: LLaMA-2-13B (left) with high Gini coefficients (concentrated outliers) uses more negative learning rates than Qwen2.5-14B (right) with low Ginis (uniform outliers). Negative LRs preserve critical, outlier-dense dimensions with high initial quality. Both plots show the gate-projection.
TRIM is designed to navigate this complexity through two primary ideas. First, it recognizes that, in many cases, reducing the variance in post-pruning quality degradation across dimensions is beneficial for overall layer stability. This is pursued via a positive learning rate in its iterative refinement process (Algorithm 1), aiming to prevent any single dimension from experiencing significant quality loss. Second, TRIM acknowledges that a simple quality equalization strategy is not universally optimal. Some output dimensions possess both high initial post-pruning quality and are critically important for the model’s function (e.g., outlier-dense dimensions as shown in Table 5). Forcing these dimensions to match the potentially lower quality of less important and more sensitive ones through aggressive variance reduction would be counterproductive. This leads to TRIM’s adaptive mechanism, which involves the monitoring of layer quality and the learning rate adaptation. When beneficial, TRIM can even employ a negative learning rate, effectively increasing quality variance. LLaMA models, for instance, often exhibit strong outlier-dense dimensions (reflected in high Gini coefficients for outlier concentration, Figure 5) and, as a result, TRIM frequently selects a negative learning rate for these models (approximately 40% of the time, compared to 10% for other models). While TRIM’s flexible learning rate and best-state selection address this inherent tension in optimization direction to a significant extent, it remains a limitation and exploring alternative adjustment mechanisms or more direct methods to balance these competing objectives could be promising for future research.
@@ 7 Conclusion
In this paper, we introduce TRIM (Targeted Row-wise Iterative Metric-driven pruning), a novel approach that, for the first time, extends targeted sparsity allocation not across layers but within each layer by introducing the concept of dimension-wise sparsity. TRIM enables stable performance at high target sparsity by iteratively refining a sparsity ratio for each output dimension of a layer. Using TRIM, we reduce the perplexity by 48% for Qwen2.5-14B and over 90% for OPT-13B compared to existing state-of-the-art methods when pruning at 80% sparsity. TRIM consistently improves pruning results when paired with layer-wise sparsity allocation techniques across OPT, LLaMA, and Qwen2.5 without extensive computational overhead. By enabling more granular control, TRIM offers a more effective and reliable path towards pruning LLMs at high sparsity targets, contributing to a deeper understanding of fine-grained sparsity allocation and paving the way for future advancements in model compression.
@@ Appendix A Quality Metric
Table 6: Perplexity results for different layer quality metrics.
Model Sparsity Cosim Flat Cosim Sample MSE
Qwen2.5-3B 60% 19.35 19.38 19.42
OPT-6.7B 70% 32.51 32.75 29.62
OPT-6.7B 80% 602.57 590.12 1970.02
Qwen2.5-14B 80% 184.36 170.23 189.66
LLaMA-2-13B 80% 162.33 173.09 142.81
@@
Table 7: Perplexity results for different dimension-wise quality metrics.
Model Sparsity Cosine Similarity PSNR MSE
Qwen2.5-3B 60% 19.35 19.33 19.26
OPT-6.7B 70% 32.51 38.64 44.72
OPT-6.7B 80% 602.57 6834.343 4405.93
Qwen2.5-14B 80% 184.36 216.87 276.01
LLaMA-2-13B 80% 162.34 226.88 137.57
@@ Appendix B Runtime Overhead
One-shot pruning methods like Wanda (Sun et al., 2024) are designed for efficient pruning of models with billions of parameters. Any extension to these methods should maintain this computational efficiency. TRIM, while iterative in nature, is designed around that and performs most computations in parallel on the GPU.
@@
Table 8: Runtime overhead of TRIM with different numbers of calibration samples.
N samples 32 64 128 256 512
Wikitext Perplexity 34.70 33.06 33.49 32.72 33.21
TRIM overhead (s) 27.76 31.78 35.96 42.45 59.88
Total Time (s) 375 562 888 1904 3325
Appendix C Robustness Analysis
@@
Appendix D Implementation Details
@@
Additionally, we give the option for input vector recalculation. One-shot pruning methods like Wanda prune block by block, and in each block layer by layer. However, when the Q,K,V-projections are pruned the activation statistic for subsequent layers (Out-projection, or the MLP block) also changes. Therefore, we include an option to recalculate the input vectors used by TRIM. Note that this does not include the Wanda metric itself, which ensures fair comparisons.
This recalculation makes pruning quality slightly better (see Table 10) and results in increased inference time. We recommend enabling this option for achieving the best results, and disabling it during hyperparameter tuning or developing.
@@
Sparsity Baseline Without Recalc With Recalc
70% 36.46 33.85 33.21
80% 348.48 186.32 180.67
Appendix E Hyperparameters

### v2

@@ 4.1 Main results
Although we primarily evaluate TRIM with Wanda (Sun et al., 2024), its dimension-wise sparsity and per-output optimization apply to any scoring rule that assigns a value to each weight. For clarity, the three metrics we test are: (i) Magnitude, which uses absolute weights $|\mathbf{W}_{ij}|$ Han et al. (2015); (ii) SparseGPT, which scores weights by the Hessian-aware ratio $\lvert\mathbf{W}\rvert^{2}/\mathrm{diag}(\mathbf{H}^{-1})]_{ij}$ with $\mathbf{H}\!\approx\!\mathbf{X}^{\top}\mathbf{X}$ from layer inputs $\mathbf{X}$ (Frantar and Alistarh, 2023); (iii) GBLM which extends Wanda by multiplying weight magnitude with an activation- and gradient-based term, $|\mathbf{W}_{ij}|\!\cdot\!(\lVert\mathbf{X}_{j}\rVert+\alpha\,\mathbf{G}_{ij})$ (i.e., Wanda + gradient score) (Das et al., 2024).
SparseGPT compares candidates over blocks of 128 inputs (columns), Magnitude compares layerwise, whereas Wanda/GBLM compare within a single output (row). We omit SparseGPT’s post-pruning weight updates to isolate the effect of pruning configuration. Table 4 shows that TRIM’s per-output configuration yields the best perplexity across all three metrics, indicating that TRIM’s gains generalize well beyond the Wanda metric.
Table 4: WikiText perplexity of LLaMA-2-13B at 70% with OWL layer ratios. TRIM improves upon the per-output configuration and generalizes to other pruning metrics. The default comparison group is underlined.
Method (layer) (128, input) (1, output) / +TRIM Magnitude 59.86 49.25 23.48 / 19.94 SparseGPT 19.75 19.23 15.19 / 14.11 GBLM 18.53 22.80 16.04 / 14.15
(a) Gini histogram.
(b) Cosine similarity vs Sparsity.
(c) LR sign vs Gini (outlier concentration).
Figure 3: Gallery of pruning diagnostics. (a) Gini histogram: Gini coefficients of the Wanda pruning metric across output dimensions. Higher Gini $\Rightarrow$ signal concentrated in fewer weights. (b) Cosine similarity vs Sparsity: Dimension-wise cosine similarities at increasing sparsity, showing heterogeneous pruning sensitivity across dimensions. (c) LR sign vs outlier Gini: LLaMA-2-13B (higher Gini; more concentrated outliers) uses more negative effective learning rates than Qwen2.5-14B (lower Gini; more uniform outliers). Both plots show the gate-projection.
@@ 5 Analysis
[…]
@@
Pruning can remove redundant weights with minimal impact at low sparsity levels. However, as sparsity rises, pruning impacts weights with essential activation signals, leading to a faster decay in performance. This non-linear decline is evident in the average trend of quality degradation on the dimension-level (Figure 3(b)) and in downstream metrics like the exponential increase in perplexity at high sparsity rates (Figure 2).
The heterogeneity of importance concentration translates directly into differing pruning sensitivities across output dimensions. To demonstrate this, we increase the uniform sparsity ratio applied to a layer and measure the resulting quality degradation for each individual output dimension. We quantify the quality using the cosine similarity $c_{i}=\text{CosineSimilarity}(Y_{i,:},\hat{Y}_{i,:})$ between the original output $Y_{i,:}$ and the pruned output $\hat{Y}_{i,:}$ for that dimension, given a set of sample inputs $X$. Figure 3(b) plots these cosine similarities $c_{i}$ against increasing levels of sparsity $T$. The rate at which quality degrades differs significantly across dimensions. Some dimensions maintain high similarity even at high sparsity levels (resilient to pruning), while others experience a rapid drop even at lower sparsity (sensitive to pruning).
Method Perplexity Baseline 8.03 $D$ w. min norm 8.19 $D$ w. max norm 273.10 Random $D$ 8.18
@@
Beyond the differences in their sensitivity to pruning, output dimensions vary significantly in their importance for model performance. To illustrate this, we fully eliminate exactly one output dimension per layer in Qwen2.5-3B, resulting in 0.03% total sparsity (Table 5). Removing the dimension with the smallest L2-norm only increases perplexity by 0.16 from the unpruned baseline, similar to choosing a random dimension (+0.15). However, pruning the dimension with the highest norm increases perplexity to 273.10.
The importance of a dimension is also linked to the concentration of outlier positions present in it. Recall that $A_{ij}$ is an outlier position if $A_{ij}>M\cdot\bar{A}$. We take the top 10% of dimensions with the largest number of outliers (i.e., "outlier-dense" dimensions) and prune them with 90% sparsity, resulting in 9% total model sparsity. Note that pruning this way retains all outlier positions since they are always the last weights to get pruned. Table 6 shows that pruning outlier-dense dimensions significantly impacts performance compared to randomly pruning an equivalent number of dimensions at the same sparsity. This result highlights two key points.
Table 6: Wikitext validation perplexity. 10% of dimensions $D$ are heavily pruned (90%). Outlier-dense dimensions are important for performance.
Model Baseline Random Outlier-dense $D$ Qwen2.5-7B 6.85 9.17 26.39 Qwen2.5-14B 5.29 7.70 26.16 LLaMA2-7B 5.47 7.27 255.06 LLaMA2-13B 4.88 6.23 74.27
First, outlier-dense dimensions are important for model performance. Second, although outlier positions in dense dimensions appear to capture most of the signal, non-outlier weights retain substantial information, limiting aggressive pruning. This effect varies across model families. For example, LLaMA models exhibit a more pronounced degradation than Qwen, indicating that outlier-dense dimensions are more important for some models than for others. Together, these findings highlight that output dimensions are not uniformly important and that simple heuristics, such as L2-Norm or outlier characteristics, can identify dimensions critical for preserving model capabilities.
@@ 6 Discussion
Targeted sparsity allocation has proven successful for enhancing the performance of pruned LLMs, but has so far only been attempted at the layer-wise level (Yin et al., 2024; Lu et al., 2024; Sun et al., 2025; Xu et al., 2025). This work extends this paradigm to individual output dimensions within each layer. Our findings show that dimension-wise sparsity is necessary, when pruning at high sparsity levels. Observation 5 demonstrates that post-pruning quality tends to degrade at an accelerated rate as sparsity increases. As a result, using coarse pruning strategies becomes increasingly detrimental. The core premise of TRIM is that significant heterogeneity exists at the dimension level (Observations 5 and 5). However, moving to such a fine-grained level of control introduces challenges: the search space for optimal sparsity configurations becomes substantially larger, and the feedback signals (e.g., quality metrics per dimension) might become less reliable or more noisy compared to layer-level aggregates.
TRIM addresses this complexity with two primary ideas. First, reducing the variance in post-pruning quality degradation across dimensions is beneficial for overall layer performance. This is pursued via a positive learning rate in its iterative refinement process, aiming to prevent any single dimension from experiencing significant quality loss. Second, TRIM recognizes that simple quality equalization is not universally optimal. Some output dimensions possess both high initial post-pruning quality and are critical to model function (e.g., outlier-dense dimensions; Table 6). Forcing these dimensions to match the potentially lower quality of less important, more sensitive ones through aggressive variance reduction would be counterproductive. This leads to TRIM’s adaptive mechanism, which involves the monitoring of layer quality and the learning rate adaptation. When beneficial, TRIM can use a negative learning rate, effectively increasing quality variance. LLaMA models, for instance, often exhibit outlier-dense dimensions (reflected in high Gini coefficients for outlier concentration, Figure 3(c)) that are important for model performance (Table 6). As a result, TRIM frequently selects a negative learning rate for these models (approximately 40% of the time, compared to 10% for other models).
@@ 7 Conclusion
In this paper, we introduce TRIM (Targeted Row-wise Iterative Metric-driven pruning), a novel approach that, extends targeted sparsity allocation not only across layers but within each layer by introducing the concept of dimension-wise sparsity. TRIM enables stable performance in challenging pruning scenarios by iteratively refining a sparsity ratio for each output. Using TRIM, we reduce perplexity by 48% for Qwen2.5-14B and over 90% for OPT-13B compared to existing state-of-the-art methods when pruning at 80% sparsity. Beyond these gains, TRIM is a lightweight, future-proof extension to importance-score pruning: it decouples scoring from allocation and accepts any importance metric as input. This plug-and-play design layers dimension-wise budgets on top of existing scorers, improving high-sparsity stability with negligible overhead, and providing a clear upgrade path as new scoring rules emerge.
Limitations
Layers with concentrated outliers benefit from negative learning rates $\alpha$, whereas more uniform layers prefer positive $\alpha$ (Fig. 3(c), Tab. 6). These findings present an inherent tension between the general benefit of reducing quality variance for overall stability and the need to preserve and potentially increase variance in highly critical dimensions. TRIM’s flexible learning rate and best-state selection address this tension, but do not eliminate it; exploring alternative update rules or explicit variance-stability trade-offs could be promising for future research.
Our allocation is unstructured within each row, and it is this flexibility that allows TRIM to outperform previous methods. However, this freedom also means it does not directly translate to fixed-sparsity, hardware-accelerated patterns such as n:m.
@@ Appendix B Runtime Overhead
One-shot pruning methods like Wanda (Sun et al., 2024) are designed for efficient pruning of models with billions of parameters. Any extension to these methods should maintain this computational efficiency. TRIM, while iterative in nature, is designed around that and performs most computations efficiently vectorized on the GPU.
@@
Table 10: Perplexity results on WikiText validation set for various models pruned at 50% and 60% sparsity. The $\Delta$ column shows the average perplexity reduction across the tested models achieved by adding TRIM. Lower is better.
Method OPT LLaMA-2 Qwen2.5 6.7B 7B 13B 3B 7B 14B $\Delta$ 0% - 10.86 5.47 4.88 8.03 6.85 5.29 - 50% Wanda 11.96 6.92 5.98 11.37 8.57 7.30 -0.94% +TRIM 11.98 6.88 5.96 11.18 8.55 7.29 60% Wanda 15.14 10.77 8.41 21.51 13.75 11.31 -11.42% +TRIM 14.68 10.64 8.26 19.06 13.05 9.98
@@
Appendix C TRIM without Layerwise Sparsity Allocation
In the main body of the paper, we evaluate TRIM in conjunction with layerwise sparsity allocation schemes such as OWL Yin et al. (2024) and AlphaPruning Lu et al. (2024). To test if the improvements of TRIM hold up without these schemes, we evaluate several models on the Wanda metric with uniform sparsity ratios (see Table 10). The improvements of TRIM without layerwise allocation schemes are comparable to the improvements with layerwise allocation schemes (compared to Table 3). In fact, since the baseline performance is typically worse without the layerwise schemes, TRIM typically demonstrates larger relative performance gains.
Appendix D Robustness Analysis
Appendix E Implementation Details
@@
Additionally, we give the option for input vector recalculation. One-shot pruning methods like Wanda prune block by block, and in each block layer by layer. However, when the Q,K,V-projections are pruned the activation statistic for subsequent layers (Out-projection, or the MLP block) also changes. Therefore, we include an option to recalculate the input vectors used by TRIM. Note that this does not include the Wanda metric itself, which ensures fair comparisons. This recalculation makes pruning quality slightly better (see Table 12) and results in increased pruning time. We recommend enabling this option for achieving the best results, and disabling it during hyperparameter tuning or developing.
@@
Sparsity Baseline Without Recalc With Recalc 70% 36.46 33.85 33.21 80% 348.48 186.32 180.67
Appendix F Generalizability of WikiText Results
We expand our perplexity evaluations to showcase additional benchmark datasets beyond WikiText, specifically C4 and the Pile, consistent with prior work. The results are summarized in Tab 13.
Model Dataset Baseline Wanda / +TRIM
Qwen2.5-7B
(65%)
WikiText 6.85 25.81 / 22.99
C4 11.88 34.64 / 31.15
Pile 7.79 22.61 / 20.39
LLaMA2-7B
(60%)
WikiText 5.47 10.77 / 10.64
C4 7.26 13.94 / 13.80
Pile 5.79 12.40 / 10.70
LLaMA2-13B
(65%)
WikiText 4.88 14.16 / 13.85
C4 6.73 19.32 / 18.98
Pile 5.44 15.42 / 15.15
Table 13: Comparison of baseline, Wanda, and Wanda+TRIM performance across different datasets.
We find that wikitext serves as a reliable indicator of overall perplexity behavior. In most cases, reductions in WikiText perplexity are accompanied by corresponding improvements on C4 and the Pile, affirming the broader effectiveness of our method beyond a single benchmark.
Appendix G Sensitivity Analysis of k
The hyperparameter k controls the number of iterations in the TRIM algorithm. A larger k allows for a more wide-ranging exploration of dimension-wise sparsity. Our findings indicate that k is a robust hyperparameter, with its impact depending on the difficulty of the pruning task. In challenging scenarios (e.g., Qwen2.5-14B at 70%), the iterative process is key, with performance improving significantly as k approaches our default of 10 before stabilizing. Conversely, in more straightforward scenarios (e.g., OPT-7B at 60%), the model’s performance is completely insensitive to k.
Table 14: Effect of different values for hyperparameter $k$ in TRIM. Perplexity is reported on Wikitext, averaged across two runs. Our default $k=10$ is underscored.
$k$ Qwen2.5-14B @ 70% OPT-7B @ 60% 3 99.46 14.73 6 66.70 14.73 10 56.67 14.73 15 58.60 14.73 25 59.98 14.73
Appendix H Hyperparameters